OpenStudy (anonymous):

evaluate the polynomial 6x-7 for 1/2n2 -1/2n and y =2

OpenStudy (anonymous):

Can't, as we don't know how any of those equations are related.

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

You haven't provided any equation that relates the three variables (x, n, y). By "relate", I mean none of the equations provided contain either of the variables in the same equation. Kind of like asking someone to convert 100 dollars to 100 yen without providing a rule for it.

OpenStudy (anonymous):

thats all it says on my paper

OpenStudy (anonymous):

Also, 1/2n2=1/4n, assuming "2n2" = 2*n*2, so 1/2n2-1/2n= -1/4n

OpenStudy (anonymous):

my answer choices are a.15 or b. -20 or c.4 or d.-16

OpenStudy (anonymous):

Have you typed the problem correctly?

OpenStudy (anonymous):